# SQ-GAN: Semantic Image Communications Using Masked Vector Quantization

### v1

@@ III-A Semantic Encoder
The encoder $E_{\mathbf{s}}$ maps the $n_{c}\times H\times W$ SSM $\mathbf{s}$ to a latent tensor $\mathbf{z}^{\mathbf{s}}=E_{\mathbf{s}}(\mathbf{s})$ with shape $C\times H_{16}\times W_{16}$, where $C=256$ is the hyperparameter representing the number of channels and $H_{16}$ and $W_{16}$ are the height and width. $E_{\mathbf{s}}$ is formed by a repeated sequence of two Residual Blocks (ResBlocks) and one down-scaling layer, that will be referred to as (ResBlock-Down). The average pooling down-scaling layer halves the height $H$ and width $W$ of the SSM. By applying 4 cascaded stages of ResBlock-Down, the final latent tensor has shape $H_{16}\times W_{16}$. The value of 4 ResBlock-Downs staged has been chosen via extensive ablation study to achieve a good balance between compression and reconstruction accuracy. In particular, changing to 3 or 5 stages has effects on the performance of the whole architecture, causing poor compression in the first case and poor semantic preservation in the second. The output of the last ResBlock-Down is processed by the final multi-head self-attention layer of the encoder as in [19].
@@ III-B Image Encoder
The image encoder $E_{\mathbf{x}}$ maps the image $\mathbf{x}$ to the latent representation $\mathbf{z}^{\mathbf{x}}$ of shape $C\times H_{16}\times W_{16}$, with conditioning provided by $\mathbf{s}$. $E_{\mathbf{x}}$ has the same structure of $E_{\mathbf{s}}$. The main difference lies in the conditioning process. In fact, not all the parts of the image $\mathbf{x}$ have the same semantic meaning and relevance: for example, in assisted driving applications, pedestrians are more important than the sky.
To help the encoder $E_{\mathbf{x}}$ to assign different importance to different parts, the image $\mathbf{x}$ is modified before encoding. The proposed method draws inspiration from the version of the Positional Embedding (PE) introduced by Dosovitskiy et al. in transformer networks for vision tasks [22]. This scheme divides the image into $16\times 16$ patches and assigns a specific PE vector to every patch in order to allow the transformer to correctly interpret the relative positions of the patches in the frame. In our case, PE as proposed in [22] is not needed since, unlike transformer architectures, the spatial correlations are preserved in convolution-based architectures. However, the idea of assigning different weights to different semantic regions of the image is used to improve the overall performance. Hence, we propose a variation of PE, referred to as SemPE. The goal of SemPE is to provide the encoder $E_{\mathbf{x}}$ with a suitable transformation of $\mathbf{x}$ and $\mathbf{s}$ that can be used to influence the feature extraction process. To this purpose, $\mathbf{s}$ is processed by SemPE, a two layer Convolutional Neural Network (CNN) designed to take into account the semantic classes of adjacent pixels. In parallel, the image $\mathbf{x}$ is processed by a one-layer CNN called $InputNet_{\mathbf{x}}$. Both outputs of SemPE and CNN are tensors of shape $128\times H\times W$. Then, as in classical PE, the input of the encoder is formed by the elementwise sum ${\rm InputNet}_{\mathbf{x}}(\mathbf{x})+{\rm SemPE}(\mathbf{s})$. This is then processed by the encoder $E_{\mathbf{x}}$ that produces the latent tensor $\mathbf{z}^{\mathbf{x}}=E_{\mathbf{x}}(\mathbf{h}_{SemPE}(\mathbf{x},\mathbf{s}))$ of shape $C\times H_{16}\times W_{16}$.
@@ III-C Semantically Conditioned Adaptive Mask Module
The latent tensors $\mathbf{z}^{\mathbf{x}}$ and $\mathbf{z}^{\mathbf{s}}$ can be interpreted as collections of $K=H_{16}W_{16}$ vectors in $\mathbb{R}^{C}$. The next step is to select the most relevant $N_{\mathbf{x}}=m_{\mathbf{x}}K$ vectors in $\mathbf{z}^{\mathbf{x}}$ and $N_{\mathbf{s}}=m_{\mathbf{s}}K$ vectors in $\mathbf{z}^{\mathbf{s}}$, where the relevance masking fraction $m_{\mathbf{x}}\in(0,1]$ and $m_{\mathbf{s}}\in(0,1]$ are design parameters. Such selection is performed by the SAMM blocks. The following discussion will consider only the latent tensor $\mathbf{z}^{\mathbf{x}}$ since the operations introduced in this section, in Section III-D and in Section III-E apply symmetrically to the $\mathbf{s}$ pipeline.
SAMM is a variation of the Adaptive Mask Module (AMM) introduced in [19]. The idea is to assign to each of the $K$ latent vectors a relevance score and then select the $N_{\mathbf{x}}$ vectors with the highest score. In the proposed SAMM, the relevance score is conditioned on the SSM $\mathbf{s}$ and the masking fraction $m_{\mathbf{x}}$ can be adjusted dynamically. This allows the network to use the same weights to compress images at different levels of compression. The differences in architecture between AMM and SAMM are shown in Fig. 3. While the classic AMM, on the left, is more suitable for general purpose applications, the SAMM on the right is designed to take into account the semantic class of the different regions of the image. The new SAMM enforces this conditioning thanks to the Spatially-Adaptive Normalization (SPADE) normalization layer [23]. SAMM takes as input the latent tensor $\mathbf{z}^{\mathbf{x}}$ and the SSM $\mathbf{s}$ and outputs a relevance score $\alpha_{k}^{\mathbf{x}}\in[0,1]$ for each latent vector $\mathbf{z}_{k}^{\mathbf{x}}$. The final step involves the multiplication of the selected latent vectors by their respective relevance scores. This is done to allow backpropagation to flow through SAMM and train its parameters, as discussed in [19].
@@ III-D Quantization and Compression
After selecting the $N_{\mathbf{x}}$ latent vectors with the highest score, the selected vectors are suitably quantized, while the non-relevant $K-N_{\mathbf{x}}$ latent vectors are just dropped. The vector quantization process follows the same steps as the classic MQ-VAE [19]. A learnable codebook $\mathcal{C}_{\mathbf{x}}=\{\mathbf{e}_{j}^{\mathbf{x}}:j=0,\ldots,J-1\}$ is used, where in this application we found convenient to use $J=1024$ codewords of dimensionality $C=256$. For each selected score-scaled relevant vector $\mathbf{z}_{k}^{{}^{\prime}\mathbf{x}}=\alpha_{k}^{\mathbf{x}}\mathbf{z}_{k}^{\mathbf{x}}$, the codeword at minimum distance is found, i.e., the codeword index $j$ is selected such that
@@
Then, the sequence of quantization indices is binary encoded using entropy coding. Every latent vector selected for quantization is located in some position in a $H_{16}\times W_{16}$ shaped array. Therefore, the quantized array is formed by $N_{\mathbf{x}}$ positions containing quantization indices (i.e., integers from 0 to $J-1$ selected as in (1)) and $K-N_{\mathbf{x}}$ “empty” positions corresponding to the discarded vectors. We assign to all discarded positions an additional special index (conventionally denoted as -1), such that we can interpret the quantized array as discrete information source over an alphabet of size $J+1=1025$. The index -1 appears with probability $1-\frac{N_{\mathbf{x}}}{K}=1-m_{\mathbf{x}}$, while the other indices from 0 to $J-1$ appear with some probabilities $\beta_{j}$ such that $\sum_{j=0}^{J-1}\beta_{j}=m_{\mathbf{x}}$. In general, this probability is unknown and depends on the specific image $\mathbf{x}$ to be encoded. However, using the fact that entropy is maximized by the uniform probability distribution, we obtain an upper bound to the length of the entropy-coded binary sequence by assuming $\beta_{j}=\frac{m_{\mathbf{x}}}{J}$. Therefore, the entropy coding rate for the index sequence is upper-bounded by:
@@
where $h_{2}(p)=-(1-p)\log_{2}(1-p)-p\log_{2}(p)$ is the binary entropy function. Since the index sequence length is $K$, the number of bits necessary to represent such a sequence is therefore upper-bounded by $B_{\mathbf{x}}=R_{\mathbf{x}}K$. Notice that this is indeed a brute-force bound, and in fact any suitable more refined compression scheme (e.g., using arithmetic coding coupled with Krichevsky-Trofimov probability sequential estimation) would achieve a lower rate. However, in this work this value is further approximated. To maintain a linear relationship between the number of bits and the number of relevant latent vectors the condition $h_{2}(p)\leq 1$ is used. By substituting $h_{2}(p)=1$ the final amount of bits is expressed as follows:
$B_{\mathbf{x}}=K+\log_{2}(J)N_{\mathbf{x}}=K(1+m_{\mathbf{x}}\log_{2}(J)).$ (3)
As will be pointed out in Section V, even under this coarse upper bound, the SQ-GAN consistently outperforms competing state-of-the-art image compression algorithms. By considering the same approach for the $N_{s}$ relevant vectors in $\mathbf{z}_{\mathbf{s}}$, $J=1024$, and normalizing by the total number of pixels $HW$, we can express the overall coding rate in BPP as a function of $m_{\mathbf{x}}$ and $m_{\mathbf{s}}$ as:  [^1] In our results, when referring to BPP, it is generally intended this total BPP unless when differently specified.
${\rm BPP}=\frac{B_{\mathbf{x}}+B_{\mathbf{s}}}{HW}=\frac{1}{256}[10(m_{\mathbf{x}}+m_{\mathbf{s}})+2].$ (4)
(notice that $K/(HW)=H_{16}W_{16}/(HW)=1/256$).
@@ III-E Tensor Reconstruction and Adaptive De-Masking Module
At the receiver, the codebooks $\mathcal{C}_{\mathbf{x}}$ and $\mathcal{C}_{\mathbf{s}}$ are stored in the decoder. The bit stream is decoded and the array of quantization indices is retrieved. Then, a tensor of shape $C\times H_{16}\times W_{16}$ is obtained by placing quantization codewords $\mathbf{e}_{j}^{\mathbf{x}}$ in correspondence of indices $j\in\{0,\ldots,J-1\}$ and a placeholder codeword denoted as $\mathbf{M}_{\mathbf{x}}$ in correspondence of all indices equal to -1 (i.e., in the position of the discarded latent space vectors). A similar operation is done for the $\mathbf{s}$ pipeline, where the placeholder is denoted by $\mathbf{M}_{\mathbf{s}}$. The placeholders are also learned in the trainign phase.
The resulting tensor is processed by the Adaptive De-Masking Module (ADM). This step is similar to what introduced in MQ-VAE [19] and makes use of a direction-constrained self-attention mechanism to gradually let the information flow from the relevant $\mathbf{e}_{j}^{\mathbf{x}}$ to the non-relevant placeholder $\mathbf{M}_{\mathbf{x}}$. This can be interpreted as a sort of non-linear “interpolation” of the latent space tensor in the discarded positions. The output $\hat{\mathbf{z}}^{\mathbf{x}}$ of the ADM is then used as the input of the decoder $D_{\mathbf{x}}$. Similarly, the output $\hat{\mathbf{z}}^{\mathbf{s}}$ of the ADM in the $\mathbf{s}$ pipeline forms the input of the decoder $D_{\mathbf{s}}$. At this point, the flows for the two pipelines for $\mathbf{x}$ and for $\mathbf{s}$ diverge again, as described in the next two subsections.
III-F Semantic Decoder
The structure of the decoder $D_{\mathbf{s}}$ is similar to a mirrored version of the encoder $E_{\mathbf{s}}$. The first input layer is composed of the multi-head self-attention layer, after which the series of s (ResBlock-Ups) is placed. Every ResBlock-Up is composed of two consecutive ResBlocks and one up-scaling layer. The up-scaling is performed by copying the value of one element in the corresponding up-scaled $2\times 2$ patch. The goal is to gradually transform the latent tensor $\hat{\mathbf{z}}^{\mathbf{s}}$ from a shape of $C\times H_{16}\times W_{16}$ back to the original shape of $n_{c}\times H\times W$, the same as the SSM $\mathbf{s}$. This decoding and up-scaling process is completed by using four consecutive ResBlock-Ups. The final reconstructed SSM $\hat{\mathbf{s}}$ is obtained by the standard practice of applying the argmax operator to assign each pixel to a specific semantic class [24]. The reconstructed SSM $\hat{\mathbf{s}}$ can now be used to condition the reconstruction $\hat{\mathbf{x}}$ of the image $\mathbf{x}$.
III-G Image Decoder
The reconstruction of $\mathbf{x}$ is obtained by applying the decoder $D_{\mathbf{x}}$ to $\hat{z}^{\mathbf{x}}$ and to the reconstructed SSM $\hat{\mathbf{s}}$. This decoder has a structure similar to the mirrored version of the encoder $E_{\mathbf{x}}$, with some key conceptual differences. The input multi-head self-attention layer has the same structure as the one used in the semantic decoder.. The sequence of four consecutive ResBlock-Up is modified to incorporate the conditioning via the SSM. The modifications focus on the normalization layers within the ResBlocks, replacing every normalization layer with the SPADE layer. This new layer is responsible for enforcing the structure of the SSM during the decoding phase [23].
IV Training and Inference
So far we have provided a structural and functional description of the blocks that form the encoder (transmitter) and decoder (receiver) in Fig. 2. This specifies the role and action of each block at runtime, but does not say how the blocks are trained. In this section, we focus on the training phase. We consider the whole proposed model in Fig. 2 as composed of two pipelines: one is used for the compression and reconstruction of $\mathbf{s}$, denoted by $G_{\mathbf{s}}$ (depicted in beige), the other for the compression and reconstruction of the $\mathbf{x}$, denoted by $G_{\mathbf{s}}$ (depicted in green). The training of these generators is obtained in an adversarial way, according to the GAN principle, i.e., the generators are coupled with corresponding discriminators, denoted as $D_{disc}^{\mathbf{s}}$ and $D_{disc}^{\mathbf{x}}$, respectively, for the sole purpose of training. The training scheme for $G_{\mathbf{s}}$ is shown in Fig. 4, and the one for $G_{\mathbf{x}}$ is shown in Fig. 5. At runtime (i.e., when the trained model is applied to a generic image for compression and reconstruction) the discriminators play no role.
@@ IV Training and Inference
In addition to the multi-step training approach, another crucial challenge is the identification and preservation of semantically relevant information. To address this problem the training process has been further improved by proposing and incorporating a data augmentation step, adding semantically relevant classes, and a semantic-aware discriminator network. The idea is to emphasize the importance of semantically relevant classes. For example, our numerical experiment we focused on dashboard camera views for assisted driving applications, where the semantic classes of interest are (for example) "traffic signs", "traffic lights", and "pedestrians". These typically occupy a small portion of the frame and are not always present in every frame. In contrast, non-relevant classes such as "sky", "vegetation", and "street", appear much more frequently and occupy large portion of the images. Therefore, data augmentation is used to prevent common but irrelevant semantic classes to dominate the training.
This section is divided as follows. Section IV-A will present the data augmentation process discussing commonly used techniques and the proposed semantic relevant classes enhancement method. Section IV-B will focus on the training of $G_{\mathbf{s}}$ while Section IV-C will focus on the training of $G_{\mathbf{x}}$ with the introduction of the proposed semantic-aware discriminator. Finally, Section IV-D will focus on the final fine-tuning process. The dataset used in training is the Cityscapes dataset [25] composed of 2975 pairs of images and associated SSMs.
IV-A Data Augmentation

### v2

@@ III-A Semantic Encoder
The encoder $E_{\mathbf{s}}$ maps the $n_{c}\times H\times W$ SSM $\mathbf{s}$ to a latent tensor $\mathbf{z}^{\mathbf{s}}=E_{\mathbf{s}}(\mathbf{s})$ with shape $C\times H_{16}\times W_{16}$, where $C=256$ is the hyperparameter representing the number of channels and $H_{16}$ and $W_{16}$ are the height and width. $E_{\mathbf{s}}$ is formed by a repeated sequence of two Residual Blocks (ResBlocks) [25] and one down-scaling layer, which will be referred to as ResBlock-Down. The average pooling down-scaling layer halves the height $H$ and width $W$ of the SSM. By applying 4 cascaded stages of ResBlock-Down, the final latent tensor has shape $H_{16}\times W_{16}$. We decided to use 4 ResBlock-Downs stages based on an extensive ablation study (not reported here because of space limitation), where this choice emerged as the best balance between compression and reconstruction accuracy. The output of the last ResBlock-Down is processed by the final multi-head self-attention layer of the encoder as in [21].
@@ III-B Image Encoder
The image encoder $E_{\mathbf{x}}$ maps the image $\mathbf{x}$ to the latent representation $\mathbf{z}^{\mathbf{x}}$ of shape $C\times H_{16}\times W_{16}$, using the conditioning provided by $\mathbf{s}$. $E_{\mathbf{x}}$ has the same structure as $E_{\mathbf{s}}$. The main difference lies in the conditioning process. In fact, not all the parts of the image $\mathbf{x}$ have the same semantic meaning and relevance: for example, in assisted driving applications, pedestrians are more important than the sky.
To help the encoder $E_{\mathbf{x}}$ assign different importance to different parts, the image $\mathbf{x}$ is modified before encoding. The proposed method draws inspiration from the version of the Positional Embedding (PE) introduced by Dosovitskiy et al. for transformer networks in vision tasks [26]. That scheme divides the image into $16\times 16$ patches and assigns a specific PE vector to each patch to allow the transformer to correctly interpret the relative positions of the patches in the frame. In our case, PE as proposed in [26] is not needed since, unlike transformer architectures, the spatial correlations are preserved in convolution-based architectures. However, the idea of assigning different weights to different semantic regions of the image is used to improve the overall performance. Hence, we propose a variation of PE, referred to as SemPE, whose goal is to provide the encoder $E_{\mathbf{x}}$ with a suitable transformation of $\mathbf{s}$ that can be used to influence the feature extraction process. SemPE processes $\mathbf{s}$ using a two layer Convolutional Neural Network (CNN) designed to take into account the semantic classes of adjacent pixels. In parallel, the image $\mathbf{x}$ is processed by a one-layer CNN called ${\rm InputNet}_{\mathbf{x}}$. The outputs of both SemPE and CNN are tensors of shape $128\times H\times W$. Then, as in the classical PE, the input of the encoder is formed by the elementwise sum ${\mathbf{h}(\mathbf{x},\mathbf{s})=\rm InputNet}_{\mathbf{x}}(\mathbf{x})+{\rm SemPE}(\mathbf{s})$. This is then processed by the encoder $E_{\mathbf{x}}$ to produce the latent tensor $\mathbf{z}^{\mathbf{x}}=E_{\mathbf{x}}(\mathbf{h}(\mathbf{x},\mathbf{s}))$ of shape $C\times H_{16}\times W_{16}$.
@@ III-C Semantically Conditioned Adaptive Mask Module
The latent tensors $\mathbf{z}^{\mathbf{x}}$ and $\mathbf{z}^{\mathbf{s}}$ can be interpreted as collections of $K=H_{16}W_{16}$ vectors of dimension $C$. The next step is to select the most relevant $N_{\mathbf{x}}=\left\lfloor m_{\mathbf{x}}K\right\rfloor$ vectors in $\mathbf{z}^{\mathbf{x}}$ and $N_{\mathbf{s}}=\left\lfloor m_{\mathbf{s}}K\right\rfloor$ vectors in $\mathbf{z}^{\mathbf{s}}$, where the relevance masking fraction $m_{\mathbf{x}}\in(0,1]$ and $m_{\mathbf{s}}\in(0,1]$ are design parameters. This selection is performed by the SAMM blocks, which we illustrate next. We only refer to the latent tensor $\mathbf{z}^{\mathbf{x}}$, as its application to the $\mathbf{s}$ pipeline is similar.
SAMM is a variation of the Adaptive Mask Module (AMM) introduced in [21]. The key idea consists of assigning to each of the $K$ latent vectors a relevance score and then select the $N_{\mathbf{x}}$ vectors with the highest score. In the proposed SAMM, the relevance score is conditioned on the SSM $\mathbf{s}$ and the masking fraction $m_{\mathbf{x}}$ can be adjusted dynamically. This allows the network to use the same weights to compress images at different levels of compression. The architecture of the conventional AMM and the new SAMM is shown in Fig. 3. While the classic AMM, on the left, is more suitable for general purpose applications, the SAMM on the right is designed to take into account the semantic class of the different regions of the image. The new SAMM enforces this conditioning thanks to the Spatially-Adaptive Normalization (SPADE) normalization layer [27]. The key advantage of this semantic conditioning is its ability to identify and prioritize the accurate representation of the regions of the image corresponding to the important semantic classes without requiring additional trainable parameters. Without the masking driven by the SSM, the encoding model would have required many more parameters to learn the features of relevant objects and correctly use this information as input of the AMM.
SAMM takes as input the latent tensor $\mathbf{z}^{\mathbf{x}}$ and the SSM $\mathbf{s}$ and outputs a relevance score $\alpha_{k}^{\mathbf{x}}\in[0,1]$ for each latent vector $\mathbf{z}_{k}^{\mathbf{x}}$. The final step involves the multiplication of the selected latent vectors by their respective relevance scores. This is done to allow backpropagation to flow through SAMM and train its parameters, as discussed in [21].
@@ III-D Quantization and Compression
After selecting the $N_{\mathbf{x}}$ latent vectors having the highest score, the selected vectors are vector-quantized, while the not relevant $K-N_{\mathbf{x}}$ latent vectors are just dropped. The vector quantization process follows the same steps as the classic MQ-VAE [21]. A learnable codebook $\mathcal{C}_{\mathbf{x}}=\{\mathbf{e}_{j}^{\mathbf{x}}:j=0,\ldots,J-1\}$ is used, where in this application we found convenient to use $J=1024$ codewords of dimension $C=256$. For each selected score-scaled relevant vector $\mathbf{z}_{k}^{{}^{\prime}\mathbf{x}}=\alpha_{k}^{\mathbf{x}}\mathbf{z}_{k}^{\mathbf{x}}$, the codeword at minimum distance is found, i.e., the codeword index $j$ is selected such that
@@
The sequence of quantization indices is then binary encoded using entropy coding. Every latent vector selected for quantization is located in some position in a $H_{16}\times W_{16}$ shaped array. Therefore, the quantized array is formed by $N_{\mathbf{x}}$ positions containing quantization indices (i.e., integers from 0 to J1J-1 selected as in (1)) and $K-N_{\mathbf{x}}$ “empty” positions corresponding to the discarded vectors. We assign to all discarded positions an additional special index (conventionally denoted as -1), so that we can interpret the quantized array as a discrete information source over an alphabet of size $J+1=1025$. The index -1 appears with probability $1-m_{\mathbf{x}}$, while the other indices from 0 to J1J-1 appear with probabilities $\beta_{j}$ such that $\sum_{j=0}^{J-1}\beta_{j}=m_{\mathbf{x}}$. In general, these probabilities are unknown and depend on the specific image $\mathbf{x}$ to be encoded. However, using the fact that entropy is maximized by the uniform probability distribution, we obtain an upper bound to the length of the entropy-coded binary sequence by assuming $\beta_{j}=\frac{m_{\mathbf{x}}}{J}$. Therefore, the entropy coding rate for the index sequence is upper-bounded by:
@@
where $h_{2}(p)=-(1-p)\log_{2}(1-p)-p\log_{2}(p)$ is the binary entropy function. Since the index sequence length is $K$, the number of bits necessary to represent such a sequence is therefore upper-bounded by $B_{\mathbf{x}}=R_{\mathbf{x}}K$. Notice that this is indeed a brute-force bound, and in fact any suitable more refined compression scheme (e.g., using arithmetic coding coupled with Krichevsky-Trofimov probability sequential estimation [28]) would achieve a lower rate. However, in this work this value is further approximated. To maintain a linear relationship between the number of bits and the number of relevant latent vectors, the condition $h_{2}(p)\leq 1$ is used. By substituting $h_{2}(p)=1$, the final number of bits is expressed as follows:
$B_{\mathbf{x}}=K(1+m_{\mathbf{x}}\log_{2}(J)).$ (3)
As will be pointed out in Section V, even under this coarse upper bound, SQ-GAN consistently outperforms competing SOTA image compression algorithms. By considering the same approach for the $N_{s}$ relevant vectors in $\mathbf{z}_{\mathbf{s}}$, $J=1024$, and normalizing by the total number of pixels $HW$, we can express the overall coding rate in BPP as a function of $m_{\mathbf{x}}$ and $m_{\mathbf{s}}$ as:  [^1] In our results, when referring to BPP, it is generally intended this total BPP, unless otherwise specified.
${\rm BPP}=\frac{B_{\mathbf{x}}+B_{\mathbf{s}}}{HW}=\frac{1}{256}[10(m_{\mathbf{x}}+m_{\mathbf{s}})+2],$ (4)
where $K/(HW)=H_{16}W_{16}/(HW)=1/256$. Notice that this formulation of the BPP considers the contribution of both the image and its SSM. Therefore, the BPP given above is fully consistent with the definition of source coding rate in image coding (coded bits per image pixel).
@@ III-E Tensor Reconstruction and Adaptive De-Masking Module
At the receiver, the codebooks $\mathcal{C}_{\mathbf{x}}$ and $\mathcal{C}_{\mathbf{s}}$ are stored in the decoder. The bit stream is decoded and the array of quantization indices is retrieved. Then, a tensor of shape $C\times H_{16}\times W_{16}$ is obtained by placing quantization codewords $\mathbf{e}_{j}^{\mathbf{x}}$ in correspondence of indices $j\in\{0,\ldots,J-1\}$ and a placeholder codeword denoted as $\mathbf{M}_{\mathbf{x}}$ in correspondence of all indices equal to -1 (i.e., in the position of the discarded latent space vectors). A similar operation is done for the $\mathbf{s}$ pipeline, where the placeholder is denoted by $\mathbf{M}_{\mathbf{s}}$. The placeholders are also learned in the training phase.
@@
The resulting tensor is processed by the Adaptive De-Masking Module (ADM). This step is similar to what was introduced in MQ-VAE [21] and makes use of a direction-constrained self-attention mechanism to gradually let the information flow from the relevant $\mathbf{e}_{j}^{\mathbf{x}}$ to the non-relevant placeholder $\mathbf{M}_{\mathbf{x}}$. This can be interpreted as a sort of non-linear “interpolation” of the latent space tensor in the discarded positions. The output $\hat{\mathbf{z}}^{\mathbf{x}}$ of the ADM is then used as the input of the decoder $D_{\mathbf{x}}$. Similarly, the output $\hat{\mathbf{z}}^{\mathbf{s}}$ of the ADM in the $\mathbf{s}$ pipeline forms the input of the decoder $D_{\mathbf{s}}$. At this point, the flows for the two pipelines for $\mathbf{x}$ and for $\mathbf{s}$ diverge again, as described in the next two subsections.
III-F Semantic Decoder
The structure of the decoder $D_{\mathbf{s}}$ is similar to a mirrored version of the encoder $E_{\mathbf{s}}$. The first input layer is composed of the multi-head self-attention layer, after which the series of ResBlock-Ups is placed. Every ResBlock-Up is composed of two consecutive ResBlocks and one up-scaling layer. The up-scaling is performed by copying the value of one element in the corresponding up-scaled $2\times 2$ patch. The goal is to gradually transform the latent tensor $\hat{\mathbf{z}}^{\mathbf{s}}$ from a shape of $C\times H_{16}\times W_{16}$ back to the original shape of $n_{c}\times H\times W$, the same as the SSM $\mathbf{s}$. This decoding and up-scaling process is completed by using four consecutive ResBlock-Ups. The final reconstructed SSM $\hat{\mathbf{s}}$ is obtained by the standard practice of applying the argmax operator to assign each pixel to a specific semantic class [29]. The reconstructed SSM $\hat{\mathbf{s}}$ can now be used to condition the reconstruction $\hat{\mathbf{x}}$ of the image $\mathbf{x}$, as described next.
III-G Image Decoder
The reconstruction of $\mathbf{x}$ is obtained by applying the decoder $D_{\mathbf{x}}$ to $\hat{z}^{\mathbf{x}}$ and to the reconstructed SSM $\hat{\mathbf{s}}$. This decoder has a structure similar to the mirrored version of the encoder $E_{\mathbf{x}}$, with some key conceptual differences. The input multi-head self-attention layer has the same structure as the one used in the semantic decoder. The sequence of four consecutive ResBlock-Up is modified to incorporate the conditioning via the SSM. The modifications focus on the normalization layers within the ResBlocks, replacing every normalization layer with the SPADE layer. This new layer is responsible for enforcing the structure of the SSM during the decoding phase [27].
IV Training and Inference
So far we have provided a structural and functional description of the blocks that compose the encoder and the decoder, as shown in Fig. 2. In this section, we illustrate the training phase of all these blocks, considering both image and SSM pipelines. We recall that the GANs rely on the interplay between a generator and a discriminator. The generator $G_{\mathbf{s}}$ of the semantic part is coupled with the corresponding discriminator, denoted as $D_{disc}^{\mathbf{s}}$, while the image generator $G_{\mathbf{x}}$ is paired to the corresponding discriminator $D_{disc}^{\mathbf{x}}$. The discriminator is a neural network that receives either an image or the SSM as input and assesses its authenticity by distinguishing between real and generated samples. It is used exclusively during training to guide the generator, providing feedback on how realistic its outputs are. This feedback enables the generator to progressively improve the quality and realism of the data it produces. The training scheme for $G_{\mathbf{s}}$ is shown in Fig. 4, while that for $G_{\mathbf{x}}$ is shown in Fig. 5. At runtime (i.e., when the trained model is applied to a generic image for compression and reconstruction), the discriminators play no role.